# Inverse problem for semidiscrete stochastic parabolic operators in arbitrary dimensions

### v1

@@ 2. Semidiscrete inverse random source problem
(3.15) $\begin{split}&J(w)+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|w|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}w|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|g|^{2}dt\\ &\leq\,C_{s_{0},\varepsilon}\left(\mathbb{E}\int_{Q}\,e^{2s\varphi}|f|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|D_{i}g|^{2}dt+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}st_{r}^{i}(e^{2s\varphi}|D_{i}w|^{2})\,dt\right.\\ &\left.+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,e^{2s\varphi}|w|^{2}\right|_{t=T}\right)+\mathcal{C}\tau^{3}e^{\tau\mathcal{C}}\|\xi\|_{H^{1}_{\mathcal{F}}(0,T;H^{1}(\partial\mathcal{M})}.\end{split}$
for all $\tau\geq\tau_{0}$, $0<h\leq h_{0}$, $sh\leq\epsilon_{0}$, where $J(w)\triangleq\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi e^{2s\theta\varphi}\,|D_{i}w|^{2}\,dt+\mathbb{E}\int_{Q}s\varphi e^{2s\theta\varphi}|A_{i}D_{i}w|^{2}\,dt$, $f\in L^{2}_{\mathcal{F}}(0,T;L_{h}^{2}(\mathcal{M}))$, $g\in L^{2}_{\mathcal{F}}(0,T;H^{1}(\mathcal{M}))$,
$\xi\in H^{1}_{\mathcal{F}}(0,T;H_{h}^{1}(\partial\mathcal{M})\cap H_{h}^{1/2}(\partial\mathcal{M)})$ satisfying compatibility condition $\xi(0)=0$, and $w$ satisfies $dw-\sum_{i\in\llbracket 1,n\rrbracket}D_{i}(\gamma_{i}D_{i}w)\,dt=fdt+gdB(t)$ with $w=\xi$ in $\partial\mathcal{M}$ and $w(0)=0$ in $\mathcal{M}$.
@@
Applying (1.10) to $z$, we have for all $\tau\geq\tau_{0}$, $0<h\leq h_{0}$, $sh\leq\epsilon_{0}$,
$\begin{split}&J(z)+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|z|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}z|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|g|^{2}dt\\ &\leq\,C_{s_{0},\varepsilon}\left(\mathbb{E}\int_{Q}\,e^{2s\varphi}|f|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|D_{i}g|^{2}dt+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}st_{r}^{i}(e^{2s\varphi}|D_{i}z|^{2})\,dt\right.\\ &\left.+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,e^{2s\varphi}|z|^{2}\right|_{t=T}\right),\end{split}$
where $J(z)\triangleq\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi e^{2s\theta\varphi}\,|D_{i}z|^{2}\,dt+\mathbb{E}\int_{Q}s\varphi e^{2s\theta\varphi}|A_{i}D_{i}z|^{2}\,dt$. Recalling that $w=z+u$ and using the triangular inequality, we can obtain that
$\begin{split}&J(w)+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|w|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}w|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|g|^{2}dt\\ &\leq\,C_{s_{0},\varepsilon}\left(\mathbb{E}\int_{Q}\,e^{2s\varphi}|f|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|D_{i}g|^{2}dt+\sum_{i=1}^{n}\mathbb{E}\int_{\Gamma^{+}}st_{r}^{i}(e^{2s\varphi}|D_{i}w|^{2})\,dt\right.\\ &\left.+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,e^{2s\varphi}|w|^{2}\right|_{t=T}+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}st_{r}^{i}(e^{2s\varphi}|D_{i}u|^{2})\,dt+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,e^{2s\varphi}|u|^{2}\right|_{t=T}\right)\\ &+J(u)+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|u|^{2}\,dt+\sum_{i,j\in\llbracket 1,n\rrbracket}\mathbb{E}\int_{Q}s^{-1}e^{2s\varphi}|D_{ij}^{2}u|^{2}\,dt.\end{split}$
Applying el Lemma 3.1 and noting that
$\begin{split}J(u)+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|u|^{2}&\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}u|^{2}\,dt+\mathbb{E}\int_{\mathcal{M}}s^{2}\,e^{2s\varphi}|u(T)|^{2}\\ &\leq\,\mathcal{C}\tau^{3}e^{\tau\mathcal{C}}\left(\|u\|_{L^{2}_{\mathcal{F}}(\Omega;L^{2}((0,T);H_{h}^{2}(\mathcal{M}))}+\|u(T)\|_{L^{2}_{\mathcal{F}_{T}}(\Omega;L_{h}^{2}(\mathcal{M})}\right),\end{split}$
we get the desired result. ∎
@@
In order to estimate the solution of (1.8) in $\mathcal{M}_{0}\times(\varepsilon,T-\varepsilon)$ by Cauchy data at $\Gamma^{+}$, we need to choose a suitable weight function $\varphi$. For this, we can choose a bounded domain $G_{1}$ with a smooth boundary such that
$G\subsetneq G_{1},\quad\overline{\Sigma}\triangleq\overline{\partial G\cap G_{1}},\quad\partial G\setminus\Sigma\subset\partial G_{1}.$
we can see that $G_{1}$ is the union of $G$ with other set $\tilde{G}$ such that $\partial\tilde{G}\cap\overline{G}=\Sigma$ and that $G_{1}\setminus\overline{G}$ have a non-empty open set.
$\begin{split}\Psi(x)>0,x\in G_{1}\quad&\Psi(x)=0,\,x\in\partial G_{1},\\ |\nabla\Psi(x)|>0,\;&x\in\overline{G_{1}}\cap\overline{G}.\end{split}$
Then, since $\overline{G}_{0}\subset G_{1}$, we can choose a sufficiently large $N>1$ such that
(3.16) $\left\{x\in G_{1}:\Psi(x)>\frac{4}{N}\|\Psi\|_{C(\overline{G_{1}})}\right\}\cap\overline{G}\supset G_{0}.$
Moreover, we choose $\beta>0$ such that
(3.17) $2\beta\varepsilon^{2}>\|\Psi\|_{C(\overline{G_{1}})}>\beta\varepsilon^{2}$
We arbitrarily fix $t_{0}\in[\sqrt{2}\varepsilon,T-\sqrt{2}\varepsilon]$. We set $\varphi(x)s(t)=\tau e^{\lambda\psi(x)-\lambda\theta(t)}$ with fixed large parameter $\lambda>0$ and $\psi(x)-\theta(t)=\Psi(x)-\beta(t-t_{0})^{2}$, $\mu_{k}=\exp{\left(\lambda\left(\frac{k}{N}\|\Psi\|_{C(\overline{G_{1}})}-\frac{\beta\varepsilon^{2}}{N}\right)\right)}$, $k=1,2,3,4$, and
(3.18) $G^{(k)}\triangleq\left\{(x,t):x\in\overline{G},\Phi(x,t)>\mu_{k}\;\text{such that}\;\Phi\triangleq e^{\lambda\psi(x)-\lambda\theta(t)}\right\}.$
In the same way as in the proof of Theorem 5.1 in Yamamoto_2009 it can be verified that
(3.19) $G_{0}\times\left(t_{0}-\frac{\varepsilon}{\sqrt{N}},t_{0}+\frac{\varepsilon}{\sqrt{N}}\right)\subset G^{(k)}\subset\overline{G}\times(t_{0}-\sqrt{2}\varepsilon,t_{0}+\sqrt{2}\varepsilon)$
@@
(3.20) $\begin{split}\partial G^{(1)}\subset\Sigma^{(1)}\cup\Sigma^{(2)},&\\ \Sigma^{(1)}\subset\Sigma\times(0,T),\quad&\quad\Sigma^{(2)}\triangleq\{(x,t):x\in G,\Phi(x,t)=\mu_{1}\}.\end{split}$
Let $\mathcal{M}^{(k)}\triangleq G^{(k)}\cap\mathcal{M}$. In order to apply Theorem 3.3, we need a cut-off function because we have no data on $\partial G^{(1)}\setminus(\Sigma\times(0,T))$. Let $\chi\in C^{\infty}(\overline{G})$ such that $0\leq\chi\leq 1$ and
(3.21) $\chi(x,t)\triangleq\left\{\begin{array}[]{cc}1&\Phi(x,t)>\mu_{3},\\ 0&\Phi(x,t)<\mu_{2}.\end{array}\right$.
@@
$\begin{split}\mathbf{F}\triangleq&\chi_{t}w-\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}(\chi)A(w))+A_{i}D_{i}(\chi)A_{i}(\gamma_{i}D_{i}w)+a_{1i}A_{i}D_{i}(\chi)A_{i}^{2}(w)\\ &+\frac{h^{2}}{4}\left(D_{i}^{2}(\chi)D_{i}(\gamma_{i}D_{i}(w))+a_{1i}D_{i}^{2}(\chi)A_{i}D_{i}(w)+2a_{1i}A_{i}D_{i}(\chi)D_{i}^{2}(w)\right).\end{split}$
By (3.20) and (3.21), we see that for all $i\in\llbracket 1,n\rrbracket$
(3.22) $v=D_{i}v=0\quad\quad\text{on}\,\,\Sigma^{(2)}\cap\partial\mathcal{M}.$
@@
$\max\left\{\Phi(x,0),\Phi(x,T)\right\}\leq\Psi(x)-2\beta\varepsilon^{2}\leq 0,$
which implies $(\mathcal{M}\times\{t=0\})\cup(\mathcal{M}\times\{t=T\})\subset\sup{(\chi)}$ and then
(3.23) $v(0)=v(T)=0.$
Applying (3.15) on $v$ and taking into account (3.22), (3.23), the choice of $\Sigma$ is arbitrary, so it is made to coincide $\Sigma\cap\partial\mathcal{M}$ with $\Gamma^{+}$ and taking $\eta_{i}$ as $t_{r}^{i}(D_{i}w)$ on $\Gamma^{+}$, we have for all $\tau\geq\tau_{0}$, $0<h\leq h_{0}$, $sh\leq\epsilon_{0}$
$\begin{split}&J(v)+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|v|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}v|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|a_{3}v|^{2}dt\\ &\leq\,C_{s_{0},\varepsilon}\left(\mathbb{E}\int_{Q}\,e^{2s\varphi}\left|\sum_{i=1}^{n}a_{1i}A_{i}D_{i}(v)+a_{2}v+\mathbf{F}\right|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|D_{i}(a_{3}v)|^{2}dt\right.\\ &\left.+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}st_{r}^{i}(e^{2s\varphi})|\eta_{i}|^{2}\,dt\right)+\mathcal{C}\tau^{3}e^{\tau\mathcal{C}}\|\xi\|^{2}_{H^{1}_{\mathcal{F}}(0,T;H_{h}^{1}(\Gamma^{+})},\end{split}$
where $J(v)\triangleq\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi e^{2s\theta\varphi}\,|D_{i}v|^{2}\,dt+\mathbb{E}\int_{Q}s\varphi e^{2s\theta\varphi}|A_{i}D_{i}v|^{2}\,dt$.
@@
(3.24) $\begin{split}J(v)+&\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|v|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}v|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|a_{3}v|^{2}dt\\ &\leq\,\mathcal{C}\tau^{3}e^{\tau\mathcal{C}}\|\xi\|^{2}_{H^{1}_{\mathcal{F}}(0,T;H_{h}^{1}(\Gamma^{+})}+C\left(\mathbb{E}\int_{Q}\,e^{2s\varphi}|\mathbf{F}|^{2}dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q}\,e^{2s\varphi}|A_{i}D_{i}v|^{2}\,dt\right.\\ &\left.+\mathbb{E}\int_{Q}e^{2s\varphi}|a_{2}v|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}\left(|A_{i}v|^{2}+|D_{i}v|^{2}\right)\,dt\right.\\ &\left.+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}st_{r}^{i}(e^{2s\varphi})|\eta_{i}|^{2}\,dt\right).\end{split}$
Noting that using (2.5), (1.12) and $e^{-2s\varphi}A_{i}(e^{2s\varphi})=1+\mathcal{O}((sh)^{2})$, for each $i\in\llbracket 1,n\rrbracket$ we have

### v2

@@ 2. Semi-discrete inverse random source problem
(3.16) $\begin{split}&J(w)+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|w|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}w|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|g|^{2}dt\\ &\leq\,\mathcal{C}_{\tau_{0},\varepsilon_{0}}\left(\mathbb{E}\int_{Q}\,e^{2s\varphi}|f|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|D_{i}g|^{2}dt+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma}st_{r}^{i}(e^{2s\varphi}|D_{i}w|^{2})\,dt\right.\\ &\left.+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,e^{2s\varphi}|w|^{2}\right|_{t=T}\right)+\mathcal{C}\tau^{3}e^{\tau\mathcal{C}}\|\xi\|^{2}_{H_{\mathcal{F}}^{1}(0,T;H_{h}^{1/2}(\partial\mathcal{M})\cap H_{h}^{1}(\partial\mathcal{M}))}.\end{split}$
for all $\tau\geq\tau_{0}$, $0<h\leq h_{0}$, $sh\leq\epsilon_{0}$, where $f\in L^{2}_{\mathcal{F}}(0,T;L_{h}^{2}(\mathcal{M}))$, $g\in L^{2}_{\mathcal{F}}(0,T;H^{1}(\mathcal{M}))$,
$\xi\in L^{2}_{\mathcal{F}}(0,T;H_{h}^{1}(\partial\mathcal{M})\cap H_{h}^{1/2}(\partial\mathcal{M)})$ satisfying compatibility condition $\left.\xi\right|_{t=0}=0$ for all $x\in\partial\mathcal{M}$ and $w$ satisfies $dw-{\sum_{i=1}^{n}}D_{i}(\gamma_{i}D_{i}w)\,dt=fdt+gdB(t)$ with $w=\xi$ on $\partial Q$ and $\left.w\right|_{t=0}=0$ in $\mathcal{M}$.
@@
Applying Theorem 3.3 to $z$, we have for all $\tau\geq\tau_{0}$, $0<h\leq h_{0}$, $sh\leq\epsilon_{0}$,
$\begin{split}&J(z)+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|z|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}z|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|g|^{2}dt\\ &\leq\,C_{\tau_{0},\varepsilon_{0}}\left(\mathbb{E}\int_{Q}\,e^{2s\varphi}|f|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|D_{i}g|^{2}dt+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma}st_{r}^{i}(e^{2s\varphi}|D_{i}z|^{2})\,dt\right.\\ &\left.+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,e^{2s\varphi}|z|^{2}\right|_{t=T}\right).\end{split}$
Recalling $w=z+u$ and using the triangular inequality, we obtain
(3.17) $\begin{split}&J(w)+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|w|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}w|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|g|^{2}dt\\ &\leq\,C_{\tau_{0},\varepsilon_{0}}\left(\mathbb{E}\int_{Q}\,e^{2s\varphi}|f|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|D_{i}g|^{2}dt+\sum_{i=1}^{n}\mathbb{E}\int_{\Gamma}st_{r}^{i}(e^{2s\varphi}|D_{i}w|^{2})\,dt\right.\\ &\left.+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,e^{2s\varphi}|w|^{2}\right|_{t=T}+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma}st_{r}^{i}(e^{2s\varphi}|D_{i}u|^{2})\,dt+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,e^{2s\varphi}|u|^{2}\right|_{t=T}\right)\\ &+J(u)+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|u|^{2}\,dt+{\sum_{i,j=1}^{n}}\mathbb{E}\int_{Q}s^{-1}e^{2s\varphi}|D_{ij}^{2}u|^{2}\,dt.\end{split}$
To estimate the term that depends on $u$ on the right-side hand of the above inequality, we note that
[…]
Now, from Lemma 3.5, we have
$\|u\|$ ${}_{L^{2}_{\mathcal{F}}(\Omega;L^{2}((0,T);H_{h}^{2}(\mathcal{M}))}+\|\left.u\right|_{t=T}\|_{L^{2}_{\mathcal{F}_{T}}(\Omega;L_{h}^{2}(\mathcal{M})}$
$\leq\mathcal{C}\left(\|u\|_{L^{2}_{\mathcal{F}}(0,T;H^{2}_{h}(\mathcal{M}))}+\|u\|_{L^{2}_{\mathcal{F}}(\Omega;C(0,T;H^{2}_{h}(\mathcal{M})))}\right)\leq\mathcal{C}\|\xi\|_{H_{\mathcal{F}}^{1}(0,T;H_{h}^{1/2}(\partial\mathcal{M})\cap H_{h}^{1}(\partial\mathcal{M}))}$
and
$\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma}t_{r}^{i}(|D_{i}u|^{2})\,dt\leq\mathcal{C}\|\xi\|^{2}_{H_{\mathcal{F}}^{1}(0,T;H_{h}^{1/2}(\partial\mathcal{M})\cap H_{h}^{1}(\partial\mathcal{M}))}.$
Combining these inequalities, we deduce that
$J(u)+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|u|^{2}\,dt+\sum_{i,j=1}^{n}$ $\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}u|^{2}\,dt+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,e^{2s\varphi}|u|^{2}\right|_{t=T}$
$+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma}st_{r}^{i}(e^{s\varphi}|D_{i}u|^{2})\,dt$ $\leq\,\mathcal{C}\|\xi\|^{2}_{H_{\mathcal{F}}^{1}(0,T;H_{h}^{1/2}(\partial\mathcal{M})\cap H_{h}^{1}(\partial\mathcal{M}))}.$
Finally, combining the above inequality with (3.17), the proof is complete. ∎
@@
In order to estimate the solution of (1.8) in $\mathcal{M}_{0}\times(\varepsilon,T-\varepsilon)$ by Cauchy data at $\Gamma\times(0,T)$. Thus, according to the geometry of $\mathcal{M}_{0}$ and $\Gamma$, we need to choose a suitable weight function $\psi$ and $\theta$, respectively.
For this, we begin by $\psi$. Choose a bounded domain $G_{1}$ with smooth boundary such that
$G\subsetneq G_{1},\quad\Gamma={\partial\mathcal{M}\cap G_{1}},\quad\partial\mathcal{M}\setminus\Gamma\subset\partial G_{1}\cap h\mathbb{Z}^{n}$
We see that $G_{1}$ is the union of $G$ with other set $\tilde{G}$ such that $\partial\tilde{G}\cap\overline{\mathcal{M}}=\Gamma$ and that $G_{1}\setminus\overline{G}$ have a non-empty open set. Choosing $\overline{\tilde{G}}_{0}\subset G_{1}\setminus\overline{G}$, we apply assumption 2.1 to obtain $\psi$ satisfying
$\psi(x)>0,x\in G_{1}$ $\psi(x)=0,\,x\in\partial G_{1},$ $|\nabla\psi(x)|>0,\quad x\in\overline{G_{1}}\cap\overline{G},$
where we see at once that $\psi=d_{0}$ also satisfies assumption 3.1 with the sets $G$ and $\Gamma$.
Since $\overline{\mathcal{M}}_{0}\subset G_{1}\cap h\mathbb{Z}^{n}$, it follows that for a sufficiently large $\tilde{N}>1$ such that
(3.18) $\mathcal{M}_{0}\subset\left\{x\in G_{1}:\psi(x)>\frac{4}{\tilde{N}}\|\psi\|_{C(\overline{G_{1}})}\right\}\cap\overline{\mathcal{M}}.$
We now turn to $\theta$. We arbitrarily fix $t_{0}\in[\sqrt{2}\varepsilon,T-\sqrt{2}\varepsilon]$ and we choose $\beta>0$ such that
(3.19) $2\beta\varepsilon^{2}>\|\psi\|_{C(\overline{G_{1}})}>\beta\varepsilon^{2}.$
Now, let us consider $\varphi=e^{\lambda\psi}$ and $\theta=e^{-\lambda\beta|t-t_{0}|^{2}}$, with fixed large parameter $\lambda>0$ and subject to the previously stated properties of $\psi$, $\beta$, and $t_{0}$, we defined the following set
[…]
with $k=1,2,3,4$. In the same way as in the proof of Theorem 5.1 in Yamamoto (2009) it can be verified that
(3.21) $\mathcal{M}_{0}\times\left(t_{0}-\frac{\varepsilon}{\sqrt{\tilde{N}}},t_{0}+\frac{\varepsilon}{\sqrt{\tilde{N}}}\right)\subset G^{(k)}\cap Q\subset\overline{\mathcal{M}}\times(t_{0}-\sqrt{2}\varepsilon,t_{0}+\sqrt{2}\varepsilon)$
@@
(3.22) $\begin{split}\partial G^{(1)}\cap\partial Q\subset\Sigma^{(1)}\cup\Sigma^{(2)},&\\ \Sigma^{(1)}\subset\Gamma\times(0,T),\quad&\quad\Sigma^{(2)}:=\{(x,t):x\in\mathcal{M},\varphi(x)\theta(t)=\mu_{1}\}.\end{split}$
In order to apply Theorem 3.7, we need a cut-off function because we have no data on $\partial G^{(1)}\setminus(\Gamma\times(0,T))$. Let $\chi\in C^{\infty}(\mathbb{R}^{n+1})$ such that $0\leq\chi\leq 1$ and
(3.23) $\chi(x,t):=\left\{\begin{array}[]{cc}1,&\varphi(x)\theta(t)>\mu_{3},\\ 0,&\varphi(x)\theta(t)<\mu_{2}.\end{array}\right$.
@@
$\begin{split}\mathbf{F}:=&\chi_{t}w-\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}(\chi)A(w))+A_{i}D_{i}(\chi)A_{i}(\gamma_{i}D_{i}w)+a_{1i}A_{i}D_{i}(\chi)A_{i}^{2}(w)\\ &+\frac{h^{2}}{4}\left(D_{i}^{2}(\chi)D_{i}(\gamma_{i}D_{i}(w))+a_{1i}D_{i}^{2}(\chi)A_{i}D_{i}(w)+2a_{1i}A_{i}D_{i}(\chi)D_{i}^{2}(w)\right).\end{split}$
By (3.22) and (3.23), we see that for all i=1,,n.i{=1,...,n.}
(3.24) $v=D_{i}v=0\quad\quad\text{on}\,\,\Sigma^{(2)}.$
@@
$\max\left\{\psi(x)-\beta|t_{0}|^{2},\psi(x)-\beta|T-t_{0}|^{2})\right\}\leq\psi(x)-2\beta\varepsilon^{2}\leq 0,$
which implies $(\mathcal{M}\times\{t=0\})\cup(\mathcal{M}\times\{t=T\})\nsubseteq\sup{(\chi)}$ and then
(3.25) $\left.v\right|_{t=0}=\left.v\right|_{t=T}=0.$
We are now in a position to use (3.16) on $v$. Taking into account (3.24), (3.25), we have for all $\tau\geq\tau_{0}$, $0<h\leq h_{0}$, $sh\leq\epsilon_{0}$
$\begin{split}&J(v)+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|v|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}v|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|a_{3}v|^{2}dt\\ &\leq\,\mathcal{C}_{\tau_{0},\varepsilon_{0}}\left(\mathbb{E}\int_{Q}\,e^{2s\varphi}\left|\sum_{i=1}^{n}a_{1i}A_{i}D_{i}(v)+a_{2}v+\mathbf{F}\right|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|D_{i}(a_{3}v)|^{2}dt\right.\\ &\left.+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma}st_{r}^{i}(e^{2s\varphi})|t_{r}^{i}(D_{i}v)|^{2}\,dt\right)+\mathcal{C}\tau^{3}e^{\tau\mathcal{C}}\|\xi\|^{2}_{H^{1}_{\mathcal{F}}(0,T;H_{h}^{1}(\Gamma))},\end{split}$
@@
(3.26) $\begin{split}J(v)+&\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|v|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}v|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|a_{3}v|^{2}dt\\ &\leq\,\mathcal{C}\tau^{3}e^{\tau\mathcal{C}}\|\xi\|^{2}_{H^{1}_{\mathcal{F}}(0,T;H_{h}^{1}(\Gamma)}+\mathcal{C}\left(\mathbb{E}\int_{Q}\,e^{2s\varphi}|\mathbf{F}|^{2}dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q}\,e^{2s\varphi}|A_{i}D_{i}v|^{2}\,dt\right.\\ &\left.+\mathbb{E}\int_{Q}e^{2s\varphi}|a_{2}v|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}\left(|A_{i}v|^{2}+|D_{i}v|^{2}\right)\,dt\right.\\ &\left.+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma}st_{r}^{i}(e^{2s\varphi})|t_{r}^{i}(D_{i}v)|^{2}\,dt\right).\end{split}$
Noting that using (2.3), (1.13) and $e^{-2s\varphi}A_{i}(e^{2s\varphi})=1+\mathcal{O}((sh)^{2})$, for each i=1,,n.i{=1,...,n.} we have